### v2

@@ 3 Quantum Reinforced Learning
Figure 2: a, the typical exploration flow for the Agent to interact with the reinforced learning objective’s target environment and obtain experience data to be stored in the replay buffer. b, the particular modular model architecture used for the Actor Model that allows the intermediate mapping $T$ that maps its input vector $\mathbf{q}_{\text{i}}$ to $\mathbf{q}_{\text{i}}$ to be implemented as 1) PQC,2)a fully connected layer,3) uniform binary number generator, and 4) constant $\mathbf{0}$ output, without changing the rest of the model design. This offers a common testbench to compare learning performance of different representative approaches.
@@
For solving RL with machine learning techniques, we can similarly identify the four processes in training: exploration, forward, backward, and update. RL problems are typically defined by its environment $Env$. An environment is fully characterized by its current state $\mathbb{s}$ and its response to an external action vector $\mathbb{a}$ at this current state. The $Env$ responses to $\mathbb{a}$ conditioned on being in $\mathbb{s}$ are the new state $\mathbb{s}^{\prime}$ of $Env$ caused by $\mathbb{a}$ and a scalar reward value, $r$, for $\mathbb{s}\xrightarrow{\mathbb{a}}\mathbb{s}^{\prime}$.
@@
Utilizing DNNs to solve an RL problem typically requires training two models together as illustated in Fig. 3. One is the Actor Model, $A(\mathbb{s})\rightarrow\mathbb{a}$, which suggests an action vector, $\mathbb{a}$, for a given observed current environment state, $\mathbb{s}$, to maximize the total reward, $R=\Sigma r$, over its future interactions with the $Env$ from current state, $\mathbb{s}$. As the $Env$ is generally not analytically predictable or even deterministic, another useful DNN-based model is the Critic Model, $C(\mathbb{s},\mathbb{a})\rightarrow R$. This model aims to self-consistently approximate the maximal total reward of an action $\mathbb{a}$ for a given environment state $\mathbb{s}$. As the Actor Model drives real actions when the trained RL model is applied to real-world open-loop problems, a more generalizable Actor Model will directly impact the real-world tasks. Therefore, we first start by building the Actor Model based on hDQNN and keep the Critical Model purely classical.
@@
After the exploration process that generates sufficient data entries in the RB, we could then sample a number ($N_{\text{b}}$) of entries randomly from the RB as a batch. As seen in Fig. 3, Elements corresponding to quantum layers’ $\mathbb{q_{\text{i}}}$ and $\mathbb{q_{\text{o}}}$ in the batched entries are extracted to train and update qtDNNs approximating quantum layers of the Actor Model to the first order within the data.
After updating qtDNNs and obtaining an updated quantum tangential approximated Actor Model $A_{\text{qt}}$ with qtDNNs as batch training compatible approximation for quantum layers, we train and update the classical Critic Model using the standard Temporal Difference (TD) learning method. For training $C$ we define a target network $C_{\text{t}}$, which is a time-delayed copy of $C$. We calculate the estimated total reward using two methods. These two methods should converge on the final values as the training for $C$ converges. Therefore, we train the prediction of $C$ consistently in the iterative training rounds. One method is straightforward, using the batched sampled values for $\mathbb{s}$ and $\mathbb{a}$ to predict total reward $R=C(\mathbb{s},\mathbb{a})$. The other method tries to predict the same value using the fact that total reward could also be calculated as $R_{\text{t}}=r+C_{\text{t}}(\mathbb{s}^{\prime},A_{\text{qt}}(\mathbb{s}^{\prime}))$ if the training for $C\simeq C_{\text{t}}$ converges to accurately predicting the total reward $R\simeq R_{\text{t}}$ and $A\simeq A_{\text{qt}}$ also suggests approximately optimal $\mathbb{a}$ that lead to the largest total future reward from $\mathbb{s}$. Training $C$ is to update its parameters such that it minimizes the loss function characterizing the difference between $R$ and $R_{\text{t}}$. Note that it is only necessary to backpropagate errors through $C$ and update its parameters this way. $C_{\text{t}}$ is updated by copying slightly older parameters of $C$ to keep a time delay between them to avoid oscillatory feedback behaviors causing instability in training. $A$ is only updated through batched backpropagation in the following step.
Now, we have updated $C$ and the qtDNNs of $A_{\text{qt}}$ with the classical DNNs shared with $A$ unchanged. In the last step, we train and update the classical DNNs of $A$ ($A_{\text{qt}}$) using the qtDNNs as an approximation for the quantum layers to the first order in the batched training. We forward pass batched $\mathbb{s}$ into $A_{\text{qt}}$ and $C$ for getting the batched predicted optimal total reward, $R=C(\mathbb{s},A_{\text{qt}}(\mathbb{s}))$. The parameters of only the classical DNNs of $A$ ($A_{\text{qt}}$) are updated to maximize the $R$ using batched backpropagation through $C$ and $A_{\text{qt}}$.
@@
The trained hDQNN-based Actor model outperforms the classical models in terms of test rewards, demonstrating a potential advantage in generalizability. The learning process with key indicators is shown in Fig.4. We noticed that both FC and 0L models underperformed PQC and FC models by a large margin in test rewards with the same exploration noise. This observation is further contrasted with the high rewards achieved in the training loop. This observation highlights entropy’s role in encouraging the model to be trained towards a generalizable global optimal in reinforced learning for complex environments [2, 3]. On the other hand, hDQNN-based model outperforms RBG-based model. This further indicates that learning a structured randomness towards the complex correlated dynamics of a complex high-dimensional environment is desired as learning decreases model entropy [2, 3] for gaining information from interacting with a previously unknown environment. We note that a simple experiment was done for the Pendulum-v1 environment provided by the standard Gymnasium library. This environment is significantly simpler with only a three-dimensional observation space and a one-dimensional action space. The results show that hDQNN-based Actor model trained with qtDNN framework performed optimally with saturating scores on par with alternatives. However, in this simpler environment, we observed that RBG-based model underperformed both quantum and classical alternatives (FC and 0L). This further highlights that even though entropy encourages generalizability in reinforced learning, unstructured randomness that can’t be trained to decrease entropy for the environment to be encoded into the model hinders learning efficiently, especially for simpler and deterministic environments.
Figure 4: The total (test) reward for the trained agents to control the humanoid to walk in Humanoid-v4 environment. The solid lines are rolling window smooth over 10 test rounds. Each test round has 10 testing epochs and each round’s averaged epoch reward is shown as dots in the plot. As the tests are carried out periodically during the training, the horizontal axis records the training epoch after which the round of tests is carried out.
Table 1: hDQNN Reinforced Learning Model Configuration for Gymnasium Humanoid-v4 Problem
@@
Components Layers Activation $N$ [^3] The number of qubits in PQC,$M$ [^4] The number of layers in PQC $d_{\text{c-link}}$ [^5] The number of classical links directly between PreDNN and PostDNN $lr$ [^6] Learning rate $\gamma$ [^7] The reward discount parameter $\tau$ [^8] Soft target model update rate
@@ 4 Discussion
During the experiments presented using hDQNN trained with the qtDNN framework in comparison with alternative classical models, we observed novel behaviors that could provide insights into the fundamental dynamics and physics in learning. These insights could be useful for implementing future improvements and applying the hDQNN-qtDNN framework to major machine learning problems including training and inferences related to the frontier large language models (LLMs).
In the experiments to understand hDQNN’s abilities to learn complex environments’ dynamics efficiently, we implemented a reinforced learning model modularity such that hDQNN’s unique PQC layer could be swapped to alternative classical mappings, which offers a direct way to compare their performances. This work demonstrated hDQNN’s potential generalizability advantage in quantum reinforced learning over classical counterparts. To the best of our knowledge, this is also the first time quantum machine learning has been applied successfully to a complex reinforcement learning task, such as solving the Humanoid-v4 problem, which could potentially be applied in advanced robotic controls.
In this work, we have provided a novel framework for efficiently training hybrid Deep Quantum Neural Networks (hDQNN) for practical, complicated tasks. We further applied this framework to reinforced learning and benchmark its performance in the standard complex Humanoid-v4 environment against comparable classical alternatives. The improved generalizability, shown through achieving the highest test rewards, of hDQNN model trained with qtDNN highlights the importance of engineered entropy in learning towards global optimal with potentially improved training data efficiency. Here, we note that the demonstrated performance of the hDQNN in our experiments for learning classical data and classical environment indicates hDQNN’s ability to model complex correlations and randomness in a complicated unknown environment. In principle, such modeling of complicated classical correlated stochastic behaviors in a classical environment could be learned purely through classical models, and there might indeed be classical models inspired by this work to solve specific reinforced learning tasks with better generalizability and data efficiency. The merit of hDQNNs trained with qtDNN is that hDQNNs have a higher degree of expressivity and could model correlated stochastic processes that are computationally hard for classical models to model [1], while it is straightforward for (noisy) quantum hardware to execute hDQNN in runtime. Training hDQNN with qtDNN framework and inference with hDQNN on (noisy) quantum hardware offer a new route in machine learning and reinforced learning research for addressing complicated real-world problems that are always stochastic, highly correlated, and even challenging to describe with low computing complexities.

### v3

@@ 3 Quantum Reinforcement Learning
Figure 2: a, the typical exploration flow for the Agent to interact with the reinforcement learning objective’s target environment, $\mathcal{E}$, and obtain experience data to be stored in the replay buffer, $\mathcal{M}$. b, the particular modular model architecture used for the Actor Model, $A$, that allows the intermediate mapping $T$ that maps its input vector $\mathbf{q}_{\text{i}}$ to $\mathbf{q}_{\text{i}}$ to be implemented as 1) PQC,2)a fully connected layer,3) uniform binary number generator, and 4) constant $\mathbf{0}$ output, without changing the rest of the model design. This offers a common testbench to compare learning performance of different representative approaches .
To build the hDQNN model for Humanoid-v4, we structured the PQC follows. The PQC is composed of $N=10$ qubits and $M=10$ layers quantum operations. For the $j$-th layer, there is, first, a group of parametrized single qubit rotations where the $i$-th qubit in the layer is rotated by the rotation operator $\mathbf{R}(\theta_{ij},\phi_{ij})$ with rotation angles provided by $2N$ output neurons of PreDNN. Following the single qubit rotations in the $j$-th layer, there is also an entanglement step that carries out one CZ-gate between $p_{j}$-th and $k_{j}$-th qubits for $p_{j}\neq k_{j}$, or reduces to identity when $p_{j}=k_{j}$, where $p_{j}$ and $k_{j}$ are produced by rounding two output neurons’ values of PreDNN. There are in total $(2N+2)M$ output neurons of PreDNN for controlling the PQC directly. The PQC’s initial state is fixed to |$\ket{0}^{\otimes N}$ and input data is encoded in the controlling parameters. $\{\theta_{ij},\phi_{ij},p_{j},k_{j}\}$.
Algorithm 1: Training an hDQNN with qtDNN for Reinforcement Learning
Input:
   Environment $\mathcal{E}$ with states $s$, actions $a$, and rewards $r$;
   Discount factor $\gamma$; maximum training iterations $N_{step}$;
   Actor model $A$ (with a Quantum Layer $\mathrm{QL}$, preDNN, and postDNN); Critic model $C$;
    following the practice of TD3, there are two Critic Models to stabilize training.
   Target (delayed) Critic model $C_{t}$ and Actor model $A_{t}$; replay buffer $\mathcal{M}$;
   Neural local approximator of QL for each sampled training data batch qtDNN;
   Learning rates for Actor (preDNN and postDNN), Critic, and qtDNN ($\alpha_{A}$, $\alpha_{C}$, $\alpha_{qt}$);
   qtDNN training times $N_{\text{qt}}$ and Actor update delays $N_{\text{A}}$ for each model update epoch.
for iteration = 1 to $N_{step}$ do
   (A) Exploration / Data Collection:
   1. Observe current state $\mathbf{s}$ from $\mathcal{E}$.
   2. Compute action $\mathbf{a}=A(\mathbf{s})$, which includes:
    (a) Forward pass through Actor’s classical layers (PreDNN) and get $\mathbf{q}_{\mathrm{i}}$.
    (b) Evaluate quantum layer $\mathrm{QL}$ on input $\mathbf{q}_{\mathrm{i}}$ to sample out $\mathbf{q}_{\mathrm{o}}$.
    (c) Pass $\mathbf{q}_{\mathrm{o}}$ and $d_{\text{c-link}}$ connections from preDNN to PostDNN to predict action $\mathbf{a}$.
      Gaussian exploration noise is added to $\mathbf{a}$
   3. Apply $\mathbf{a}$ in environment $\mathcal{E}$, observe reward $r$ and next state $\mathbf{s}^{\prime}$.
   4. Store transition $(\mathbf{s},\mathbf{a},\mathbf{q}_{\mathrm{i}},\mathbf{q}_{\mathrm{o}},r,\mathbf{s}^{\prime})$ in replay buffer $\mathcal{M}$.
   (B) Update DNNs:
   Sample a batch $\mathcal{B}$ containing $N_{\text{b}}$ entries of $(\mathbf{s},\mathbf{a},\mathbf{q}_{\mathrm{i}},\mathbf{q}_{\mathrm{o}},r,\mathbf{s}^{\prime})$ from $\mathcal{M}$
    (B1) Tangential Network (qtDNN) Update:
    1. Sample $N_{\text{qt}}$ tiny-batches $\mathcal{B}_{\text{qt}}$} from $\mathcal{B}$ containing $\bigl{(}\mathbf{q}_{\mathrm{i}},\mathbf{q}_{\mathrm{o}}\bigr{)}$}.
    2. for $\mathcal{B}_{\text{qt}}$ in $\mathcal{B}_{\text{qt}}$} do
      Update qtDNN’s parameters $\theta_{\text{qt}}$ (with learning rate $\alpha_{\text{qt}}$) by minimizing
       $\mathrm{BCE\,Loss}(\text{qtDNN}(\mathbf{q}_{\mathrm{i}}),\mathbf{q}_{\mathrm{o}})$ for $\mathcal{B}_{\text{qt}}$.
     end for
    (B2) Critic Update:
    1. For each sample, compute next-action $\mathbf{a}_{\text{t}}=A_{\text{t,qt}}(\mathbf{s}^{\prime})$,
     where $A_{\text{t,qt}}$ is the surrogate target Actor $A_{\text{t}}$
    2. Compute Critic target:
     $R_{\text{t}}=r+\gamma\,C_{\text{t}}\bigl{(}\mathbf{s}^{\prime},\hat{\mathbf{a}}\bigr{)}.$
    3. Update $C$’s parameters $\theta_{C}$ by minimizing the TD-loss $\|C(\mathbf{s},\mathbf{a})-R_{\text{t}}\|$.
    (B3) Actor Update (via qtDNN) if iteration mod $N_{\text{A}}$ is 0 :
    1. Compute actions $\hat{\mathbf{a}}=A_{\text{qt}}(\mathbf{s})$ for each $\mathbf{s}$.
     where $A_{\text{qt}}$ is the Actor $A$ with the quantum layer replaced by qtDNN.
    2. Update Actor’s classical parameters, $\theta_{A}$ by maximizing Critic’s value:
     $R=C\bigl{(}\mathbf{s},A_{\text{qt}}(\mathbf{s})\bigr{)}$
     via standard backprop through qtDNN instead of QL.
    (B4) Target Critic Update if iteration mod $N_{\text{A}}$ is 0 :
     Soft update $\theta_{C_{\text{t}}}\leftarrow\tau\theta_{C}+(1-\tau)\theta_{C_{\text{t}}}$, where $10<\tau<$ 1.
     Soft update $\theta_{A_{\text{t}}}\leftarrow\tau\theta_{A}+(1-\tau)\theta_{A_{\text{t}}}$, where $10<\tau<$ 1.
end for
Output: Final trained Actor $A$ (with quantum layer $\mathrm{QL}$) and Critic $C$.
Table 1: Pseudocode for training a hybrid Deep Quantum Neural Network (hDQNN) with qtDNN-based tangential approximation in a reinforcement learning setting.
@@ 2. Update Actor’s classical parameters, $\theta_{A}$ by maximizing Critic’s value:
For solving RL with machine learning techniques, we can similarly identify the four processes in training: exploration, forward, backward, and update. RL problems are typically defined by its environment $\mathcal{E}$. An environment is fully characterized by its current state $\mathbf{s}$ and its response to an external action vector $\mathbf{a}$ at this current state. The $\mathcal{E}$ responses to $\mathbf{a}$ conditioned on being in $\mathbf{s}$ are the new state $\mathbf{s}^{\prime}$ of $\mathcal{E}$ caused by $\mathbf{a}$ and a scalar reward value, $r$, for $\mathbf{s}\xrightarrow{\mathbf{a}}\mathbf{s}^{\prime}$.
@@
Utilizing DNNs to solve an RL problem typically requires training two models together as illustrated in Fig. 3. One is the Actor Model, $A(\mathbf{s})\rightarrow\mathbf{a}$, which suggests an action vector, $\mathbf{a}$, for a given observed current environment state, $\mathbf{s}$, to maximize the total reward, $R=\Sigma r$, over its future interactions with the $\mathcal{E}$ from current state, $\mathbf{s}$. As the $\mathcal{E}$ is generally not analytically predictable or even deterministic, another useful DNN-based model is the Critic Model, $C(\mathbf{s},\mathbf{a})\rightarrow R$. This model aims to self-consistently approximate the maximal total reward of an action $\mathbf{a}$ for a given environment state $\mathbf{s}$. As the Actor Model drives real actions when the trained RL model is applied to real-world open-loop problems, a more generalizable Actor Model will directly impact the real-world tasks. Therefore, we first start by building the Actor Model based on hDQNN and keep the Critical Model purely classical.
@@
After the exploration process that generates sufficient data entries in the $\mathcal{M}$, we could then sample a number ($N_{\text{b}}$) of entries randomly from the $\mathcal{M}$ as a batch. As seen in Fig. 3, Elements corresponding to quantum layers’ $\mathbf{q_{\text{i}}}$ and $\mathbf{q_{\text{o}}}$ in the batched entries are extracted to train and update qtDNNs approximating quantum layers of the Actor Model to the first order within the data.
After updating qtDNNs and obtaining an updated surrogate Actor Model $A_{\text{qt}}$ with qtDNNs replacing PQCs of $A$ as batch training compatible moving local approximation for quantum layers, we train and update the classical Critic Model using the standard Temporal Difference (TD) learning method. For training $C$ ($A$) we define a target network $C_{\text{t}}$ ($A_{\text{t}}$), which is a time-delayed copy of $C$ ($A$). Similarly, we also define surrogate target network $A_{\text{t,qt}}$ from $A_{\text{qt}}$ with its PQCs similarly replaced with qtDNNs. We calculate the estimated total reward using two methods. These two methods should converge on the final values as the training for $C$ converges. Therefore, we train the prediction of $C$ consistently in the iterative training rounds. One method is straightforward, using the batched sampled values for $\mathbf{s}$ and $\mathbf{a}$ to predict total reward $R=C(\mathbf{s},\mathbf{a})$. The other method tries to predict the same value using the fact that total reward could also be calculated as $R_{\text{t}}=r+C_{\text{t}}(\mathbf{s}^{\prime},A_{\text{t,qt}}(\mathbf{s}^{\prime}))$ if the training for $C\simeq C_{\text{t}}$ converges to accurately predicting the total reward $R\simeq R_{\text{t}}$ and $A\simeq A_{\text{qt}}$ also suggests approximately optimal $\mathbf{a}$ that lead to the largest total future reward from $\mathbf{s}$. Training $C$ is to update its parameters such that it minimizes the loss function characterizing the difference between $R$ and $R_{\text{t}}$. Note that it is only necessary to backpropagate errors through $C$ and update its parameters this way. $C_{\text{t}}$ is updated by copying slightly older parameters of $C$ to keep a time delay between them to avoid oscillatory feedback behaviors causing instability in training. $A$ is only updated through batched backpropagation in the following step.
Now, we have updated $C$ and the qtDNNs of $A_{\text{qt}}$ with the classical DNNs shared with $A$ unchanged. In the last step, we train and update the classical DNNs of $A$ ($A_{\text{qt}}$) using the qtDNNs as an approximation for the quantum layers to the first order in the batched training. We forward pass batched $\mathbf{s}$ into $A_{\text{qt}}$ and $C$ for getting the batched predicted optimal total reward, $R=C(\mathbf{s},A_{\text{qt}}(\mathbf{s}))$. The parameters of only the classical DNNs of $A$ ($A_{\text{qt}}$) are updated to maximize the $R$ using batched backpropagation through $C$ and $A_{\text{qt}}$.
@@
The trained hDQNN-based Actor model outperforms the classical models in terms of test rewards, demonstrating a potential advantage in generalizability. The learning process with key indicators is shown in Fig.4. We noticed that both FC and 0L models underperformed PQC and FC models by a large margin in test rewards with the same exploration noise. This observation is further contrasted with the high rewards achieved in the training loop. It highlights entropy’s role in encouraging the model to be trained towards a generalizable global optimal in reinforcement learning for complex environments [5, 6]. On the other hand, hDQNN-based model outperforms RBG-based model. It further indicates that learning a structured randomness towards the complex correlated dynamics of a complex high-dimensional environment is desired as learning decreases model entropy [5, 6] for gaining information from interacting with a previously unknown environment. We note that a simple experiment was done for the Pendulum-v1 environment provided by the standard Gymnasium library. The environment is significantly simpler with only a three-dimensional observation space and a one-dimensional action space. The results show that hDQNN-based Actor model trained with qtDNN framework performed optimally with saturating scores on par with alternatives. However, in this simpler environment, we observed that RBG-based model underperformed both quantum and classical alternatives (FC and 0L). The result further highlights that even though entropy encourages generalizability in reinforcement learning [6], unstructured randomness that can’t be trained to decrease entropy for the environment to be encoded into the model hinders learning efficiently, especially for simpler and deterministic environments.
Figure 4: The total test reward for the trained agents to control the humanoid to walk in Humanoid-v4 environment. The solid lines are rolling window smooth over 10 test rounds. Each test round has 10 testing episode and each round’s averaged episode reward is shown as dots in the plot. As the tests are carried out periodically during the training, the horizontal axis records the training episode after which the round of tests is carried out. The dashed lines correspond to performance of classical RL models implemented with widely used SOTA architectures [17, 18, 19].
Table 2: hDQNN Reinforcement Learning Model Configuration for Gymnasium Humanoid-v4 Problem
@@
Components Layers Activation $N$ [^3] The number of qubits in PQC,$M$ [^4] The number of layers in PQC $d_{\text{c-link}}$ [^5] The number of classical links directly between PreDNN and PostDNN $\alpha$ [^6] Learning rate $\gamma$ [^7] The reward discount parameter $\tau$ [^8] Soft target model update rate
@@ 4 Discussion
Comparing the hDQNN trained with the qtDNN framework with alternative classical models, we observed novel behaviors that could provide insights into the fundamental dynamics and physics in learning. These insights could be useful for implementing future improvements and applying the hDQNN+qtDNN framework to major machine learning problems including training and inferences related to the frontier large language models (LLMs) and advanced robotics.
In the experiments to understand hDQNN’s abilities to learn a complex environment’s dynamics efficiently, we implemented a hDQNN reinforcement learning model where its PQC layer could be swapped to alternative classical mappings. This offers a direct way to compare PQC and classical alternatives with other variables kept the same. This work demonstrated hDQNN’s potential generalizability advantage in quantum reinforcement learning over classical counterparts including public models implemented with SOTA architectures. To the best of our knowledge, this is also the first time quantum machine learning has been applied successfully to a complex high-dimensional reinforcement learning task like the Humanoid-v4, which could potentially be applied in advanced robotics.
In this work, we have provided a novel architecture for efficiently training hDQNN for practical, complicated ML tasks. We further applied this architecture to reinforcement learning and benchmark its performance in the complex high-dimensional Humanoid-v4 environment against classical alternatives. The improved generalizability, shown through achieving the highest test episode rewards (see Fig. 4), of hDQNN model trained with qtDNN highlights the importance of learning stochastic correlations and optimizing entropy in challenging reinforcement learning tasks. In principle, modeling of complicated classical correlated stochastic behaviors in a classical environment could be learned purely through classical models, and there might indeed be classical models inspired by this work to solve specific reinforcement learning tasks. The merit of hDQNNs trained with qtDNN is that hDQNNs have a higher degree of expressivity, while being straightforward for (noisy) hybrid quantum classical hardware to execute hDQNN in runtime, could model correlated stochastic processes that are hard for classical models to simulate [8] with comparable classical resources. Training hDQNN+qtDNN and inferencing with hDQNN on (noisy) quantum hardware offer a new route in reinforcement learning and AI research for addressing complicated real-world problems that are always stochastic, highly correlated, and even challenging to describe with low computing complexities.